OpenStudy (anonymous):

For an arithmetic series, given s12=186 and t20=83, what is s40?

OpenStudy (amistre64):

\[\frac{12(t_1+t_{12})}{2}=186\] \[t_{20} = 83\] \[\frac{40(t_1+t_{40})}{2}=\_\_\]

OpenStudy (amistre64):

or do we start at a t0?

OpenStudy (amistre64):

186/6 = 31 - t12 = t1 t1 + d(19) = 83 31 - t12 + d(19) = 83 - t12 + d(19) = 52 t12 = 52 - d(19) hmm

OpenStudy (amistre64):

got a bad negative :) t12 = -52 + d(19)

OpenStudy (amistre64):

should we start at t0? the mean of t0 and t40 is t20 then

OpenStudy (amistre64):

but thats 41 terms ... not 40

OpenStudy (anonymous):

sorry i have no idea but thx for trying to help :)

OpenStudy (amistre64):

\[2t_1+11d=31\]\[t_1+19d=83\] \[d = \frac{-2(83)+31}{-2(19)+11}=5\] \[t_1=83-19(5)=-12\] \[t_{40}=-12+5(39)=183\] \[20(-12+183)=3420\]

OpenStudy (amistre64):

somethings off just a tad

OpenStudy (amistre64):

nah, it should be fine

OpenStudy (anonymous):

yeah thats the answer thanks!